OpenStudy (idily101):

my other question is similar 24y-2(6-y)=6(3y+2)

OpenStudy (idily101):

so far I got 24y-12y-2y=18y+12 is this right so far?

OpenStudy (jim766):

-2 * 6 is not 12y

OpenStudy (jim766):

the rest looks good

OpenStudy (idily101):

oh it should be -12y?

OpenStudy (jim766):

24y - 12 -2y = 18y + 12 simplify each side first

OpenStudy (idily101):

then subtract 12 from both sides

OpenStudy (jim766):

opps and -2 * -y = 2y 24y - 12 + 2y = 18y + 12

OpenStudy (jim766):

before that I would combine 24y and 2y on the left side

OpenStudy (idily101):

oh to equal 26y

OpenStudy (jim766):

right 26y - 12 = 18y + 12 now I would get rid of the smallest variable

OpenStudy (idily101):

then you would subtract 12?

OpenStudy (idily101):

or add 12

OpenStudy (jim766):

you could, but do you see how there is a variable on both sides?

OpenStudy (jim766):

I like to fix that first, you are trying to get the variable alone right?

OpenStudy (jim766):

but if you have a variable on both side, that is hard to do....so how would you get rid of the 18y? what is the opposite of + 18y

OpenStudy (idily101):

-18y

OpenStudy (jim766):

right! 26y - 12 = 18y + 12 -18y -18y

OpenStudy (jim766):

what is 26y - 18y

OpenStudy (idily101):

-8y on the left side

OpenStudy (jim766):

Good! 26y - 12 = 18y + 12 -18y -18y 8y -12 = 12 now get the y by itself what do you want to do?

OpenStudy (idily101):

divide by 8

OpenStudy (jim766):

you are going to need to divide by 8, but not yet... first get rid of the -12, what is the opposite of -12

OpenStudy (idily101):

you add im sorry algebra is sorta for me glad that you took ur time to help me

OpenStudy (jim766):

np right 8y -12 = 12 +12 =12 8y = 24 , now divide by 8 on both sides

OpenStudy (idily101):

to equal y=3

OpenStudy (jim766):

right! good job

OpenStudy (idily101):

thanks jim appericiate the help from u therefore ur my fan

OpenStudy (jim766):

yw anytime